Suppose the demand function for a firm’s product is given by 𝐿𝑛 𝑄π‘₯ 𝑑 = 7 βˆ’ 1.5𝐿𝑛𝑃π‘₯ + 2𝐿𝑛𝑃𝑦 βˆ’ 0.5𝐿𝑛𝑀 + 𝐿𝑛𝐴 Where Px = $15, Py = $6, M = $40,000, and A =$350. a. Determine the own price elasticity of demand, and state whether demand is elastic, inelastic, or unitary elastic. (3mks) b. Determine the cross-price elasticity of demand between good X and good Y, and state whether these two goods are substitutes or complements. (3mks) c. Determine the income elasticity of demand, and state whether good X is a normal or inferior good. (3mks) d. Determine the own advertising elasticity of demand. (3mks


1
Expert's answer
2022-01-31T09:57:20-0500

"lnQ_xd=7-1.5lnP_x+2lnP_y-0.5lnM+lnA"

"P_x=" $"15" , "P_y=" $"6" , "M=40000, A=" $"350"

"lnQ=7-1.5ln(15)+2ln(6)-0.5ln(40000)+ln(350)"

"lnQ=7-4.0624+3.5835-5.2983+5.8579"

"lnQ=7.0811"

"Q=e^{7.0811}=1189.276"

a)"E_p=\frac{dQ}{dP}\times \frac{P}{Q}=-1.5\times\frac{350}{1189.276}=-0.441, inelastic""demand"

b)"E_cpx=" "\frac{dQ}{dP_X}\times \frac{P_X}{Q}, P_X="$"15"

"=-1.5\times\frac{15}{1189.276}=-0.0189"

"E_cpx=\frac{dQ}{dP_Y}\times \frac{P_Y}{Q},P_Y=" $"6"

"=2\times\frac{6}{1189.276}=0.01"

X and Y are substitutes.

c)"E_I=\frac{dQ}{dI}\times \frac{I}{X}=0.5\times \frac{40000}{15}=-1.333"

X is an inferior good.

d) "AED=\frac{dQ}{dA}\times \frac{A}{Q}=1\times\frac{350}{1189.276}=0.29"